# LookAhead: Preventing DeFi Attacks via Unveiling Adversarial Contracts

### v4

@@ Abstract.
Decentralized Finance (DeFi) incidents stemming from the exploitation of smart contract vulnerabilities have culminated in financial damages exceeding $3$ billion US dollars. Existing defense mechanisms typically focus on detecting and reacting to malicious transactions executed by attackers that target victim contracts. However, with the emergence of private transaction pools where transactions are sent directly to miners without first appearing in public mempools, current detection tools face significant challenges in identifying attack activities effectively.
Based on the fact that most attack logic rely on deploying one or more intermediate smart contracts as supporting components to the exploitation of victim contracts, in this paper, we propose a new direction for detecting DeFi attacks that focuses on identifying adversarial contracts instead of adversarial transactions. Our approach allows us to leverage common attack patterns, code semantics and intrinsic characteristics found in malicious smart contracts to build the LookAhead system based on Machine Learning (ML) classifiers and a transformer model that is able to effectively distinguish adversarial contracts from benign ones, and make just-in-time predictions of potential zero-day attacks. Our contributions are three-fold: First, we construct a comprehensive dataset consisting of features extracted and constructed from $210,643$ recent contracts and $375$ hand-picked then carefully reviewed and identified adversarial contracts deployed on the Ethereum and Binance Smart Chain (BSC) blockchains. Secondly, we design a condensed representation of smart contract programs called Pruned Semantic-Control Flow Tokenization (PSCFT) and use it to train a combination of ML models that understand the behaviour of malicious codes based on function calls, control flows and other pattern-conforming features. Lastly, we provide the complete implementation of LookAhead and the evaluation of its performance metrics for detecting adversarial contracts. Our experiments show that LookAhead achieves an F1-score as high as $0.8966$, which represents an improvement of over $44.4\%$ compared to the previous state-of-the-art solution Forta, with a False Positive Rate (FPR) at only $0.16\%$.
copyright: none
@@ 1. Introduction
Decentralized Finance (DeFi) has gained significant traction within the blockchain ecosystem over the recent years and is continuing to evolve rapidly, stimulating the rise of a diverse array of DeFi applications such as: decentralized exchanges (DEXs) (Adams et al., [n. d.]; Martinelli and Mushegian, [n. d.]), lending platforms (AAV, 2024; Robert and Geoffrey, [n. d.]), and derivatives instruments (Juliano, [n. d.]). The total value locked (TVL) in these DeFi protocols has exceeded \$100 billion in mid 2024 (Def, 2024). Among the various blockchain networks, Ethereum (Wood et al., 2014) and Ethereum Virtual Machine (EVM)-compatible chains (Arb, 2024; BNB, 2024; Opt, 2024; Pol, 2024) have formed a particularly thriving DeFi ecosystem, collectively they play a pivotal role in shaping the future of the blockchain landscape. At the heart of Decentralized Finance are smart contracts, which are self-executing programs deployed on blockchains, enabling trustless and transparent financial interactions without the need for intermediaries. The functionalities of smart contracts are invoked by users through transactions, after which the blockchain consensus algorithms ensure predefined actions and state transitions are execute correctly and autonomously on-chain. However, just like traditional software, smart contracts are fundamentally code-based programs, which are susceptible to vulnerabilities and various other security risks. In fact, DeFi incidents resulting from smart contract attacks have caused over 3 billion US dollars in financial losses (DeF, 2024; Lea, 2024; Zhou et al., 2023).
@@
Limitations of Existing Work. Existing real-time defense mechanisms leverage heuristics (Qin et al., 2023b; Wu et al., 2023; Zhang et al., 2020; Zhou et al., 2020) and Machine Learning (ML) (Gai et al., 2023; Wang et al., 2022) techniques to detect and respond to yet-to-be-confirmed transactions in public mempools that are adversarial in nature. The main issue with this transaction-based detection approach is that it faces significant challenges in identifying private adversarial transactions. Namely, attackers can utilize private mempool services to send adversarial transactions directly to miners, evading visibility from other participants of the blockchain network before the transaction is confirmed and effectively bypassing attack detection. The ability to send private transactions was originally developed by researchers and industry practitioners to counter malicious activities on-chain such as front-running, but such services have unfortunately been increasingly abused by attackers for vulnerability exploitation purposes.
Figure 1. Percentage of DeFi attacks using
private transactions between 2022 and 2023.
Figure 2. Cumulative distribution function (CDF) of the rescue time frame in DeFi attacks.
Key Observations. Based on our empirical study of the historical DeFi incidents over the past few years, we found that out of the $142$ DeFi attacks on Ethereum collected by us, $80$ cases ($56.3\%$) involve transactions using private mempool services. Moreover, as shown in Figure 2, the proportion of adversarial transactions sent using private services exhibits a significant upward trend, rising from $44.4\%$ in the first half of 2022 to a noticeably higher $77.8\%$ in the latter half of 2023. We emphasize the critical importance of establishing a new approach for detecting DeFi attacks that can remain effective even when private mempool services are employed by attackers. Another important observation made by (Zhou et al., 2023) is that most attacks are not executed atomically within the constructor during adversarial contract deployment, which provides a detection and rescue time frame for defenders and potential victims. As illustrated in Figure 2, over $60\%$ of attacks exhibit a difference between the confirmation time of $tx_{deploy}$ and $tx_{first}$ that satisfies: $t_{first}-t_{deploy}\geq 100s$. Based on the aforementioned studies and observations, we design our method for preventing DeFi attacks via unveiling adversarial contracts.
Our method. In this paper, we propose the LookAhead system, representing a new direction for detecting DeFi attacks, i.e., focusing on identifying adversarial contracts instead of adversarial transactions, and allowing us to proactively identify potential attack intentions and victims, even in the case of private adversarial transactions.
@@
By shifting the focus to analyzing and identifying adversarial contracts, we filled several gaps currently present in traditional transaction-based detection tools. We first make an empirical observation that adversarial contracts used in DeFi attacks often contain similar patterns, for example, over 70% of the contracts in our adversarial contract dataset use funds from anonymous sources for deployment; more than 98% are closed-sourced, etc. We also note that depending on the vulnerabilities targeted, attackers write contract code in specific manners. For instance, flashloan and reentrancy attacks are two of the most notorious and frequently studied exploitation. Both of them require the establishment of complex function call chains and the implementation of attack logic within callback functions. To use these behaviours to our advantage for identifying adversarial contracts, we using the stacking technique to design three ML models, one transformer, multiple candidate classifiers, and a meta classifier that, when used together, can understand the intrinsic characteristics of malicious smart contracts based on code semantics and common patterns. To support the ML models in understanding the behaviour of Solidity smart contract code, we develop a specialized tokenization method called Pruned Semantic-Control Flow Tokenization (PSCFT) that is integrated directly into a custom contract bytecode decompiler. By leveraging modern ML architecture and models, we can produce batch predictions for newly created contracts on-chain in a timely manner with amortized $t^{pred}\ll 100s$. The result of our design is LookAhead, a system that is able to effectively distinguish adversarial contracts from benign ones, and make just-in-time predictions of potential zero-day attacks.
@@
To ensure the performance and generality of our ML models, we hand-picked then carefully reviewed and identified $375$ adversarial contracts used in DeFi incidents between April 2020 and July 2024 from both the Ethereum and Binance Smart Chain (BSC) to form our attack contract dataset. And in total $796,437$ contracts deployed on Ethereum from June 2022 to June 2024 were collected, with $210,643$ of them being used as benign samples after filtration. We extract a selection of features from the contracts capturing patterns exhibited during both their implementation and their deployment stages. Given that our dataset contains a combination of contracts from both recent and historical attacks, we believe that our feature selection will remain reasonably relevant and effective even in the rapid growing landscape of decentralized finance.
@@
We build our classifiers using multiple supervised ML algorithms and conduct a comprehensive performance evaluation. Experiment results show that our method for identifying adversarial contracts performs exceptionally well with an F1-Score produced by the KNN meta classifier reaching as high as $0.8966$, and a false positive rate of only $0.16\%$, which represents an improvement of over $44.4\%$ compared to the previous state-of-the-art solution Forta.
@@
  • We propose a new direction for detecting DeFi attacks that focuses on identifying adversarial contracts instead of adversarial transactions.
  • A comprehensive dataset of adversarial and benign smart contracts with a variety of features pre-extracted, pre-processed and attached to each item.
  • A smart contract bytecode lifting and analysis pipeline with an integrated protocol to extract and generate features and PSCFT for transformer and classifier training and fine-tuning.
  • An ML-based system consisting of two open-weights classifiers and one meta classifier model that is pre-trained, fine-tunable, and capable of conducting accurate and just-in-time detection of potential adversarial attacks, even when private adversarial transactions are employed by attackers.
  • The implementation and a detailed evaluation as well as performance analysis of LookAhead’s detection system and methods.
@@ 2.1. Ethereum & Decentralized Finance (DeFi)
Blockchain is a distributed ledger built on top of a peer-to-peer (P2P) network, maintained through consensus mechanisms to ensure the consistency of the ledger. Ethereum (Wood et al., 2014) is one of the most used blockchain platforms. It was the first to introduce the Ethereum Virtual Machine (EVM) that support smart contracts, and has inspired the development of a series of EVM-compatible chains. DeFi is a blockchain-based financial ecosystem powered by smart contracts to offer financial services in a more open and transparent manner. There are several unique concepts in DeFi that differ from traditional finance.
@@
Transactions. Transactions are used to transfer Ether, invoke contracts, or create contracts. After a transaction is sent, it is first placed in the mempool of a blockchain node. Public mempools are transparent and openly accessible and filled with pending transactions, waiting to be selected and finalized by block producers.
@@ 2.2. Maximal Extractable Value (MEV)
Traditionally, miner determine the order of transactions in a block based on gas prices. Since the mempool is publicly accessible, users have the ability to influence the placement of their transactions by adjusting the gas price, allowing for front-running (Eskandari et al., 2020; Torres et al., 2021) or back-running to happen. The value derived from manipulating the order of transactions within a block, beyond the standard block rewards and transaction fees, is referred to as Maximal Extractable Value (MEV) (Daian et al., 2020). It has been proven to have a significant impact on the security of blockchain networks (Qin et al., 2022).
Front-running. Front-running is a fundamental means used to extract MEV, typically employed in arbitrage, liquidation, and sandwich attacks. Front-runners monitor the public mempool for victim transactions that meets specific conditions, and attempt to raise the gas price to execute their own transactions first against the original transaction. Furthermore, front-running has also been used in intrusion prevention systems (Qin et al., 2023a; Zhang et al., 2023), as a way to front-run adversarial transactions in order to prevent DeFi attacks.
Private Mempool Services. In response to MEV, researchers and industry practitioners have proposed various countermeasures (Yang et al., 2022). The MEV-auction solution introduced by Flashbots (Fla, 2024) has recently gained popularity (Weintraub et al., 2022). This solution introduced the concept of relay systems (Lyu et al., 2022), which provide private channels for transactions to be submitted directly from the users to block producers without being broadcasted to the network, ensuring that transactions remain private until they are finalized, preventing them from being monitored or intercepted by others. However, such private mempool service are also utilized by attackers to send private adversarial transactions, aiming to execute attack logic and effectively bypass various detection methods.
@@ 3.1. Motivating Example
On July 21, 2023, DeFi protocol Conic Finance suffered a major exploit (DLNews, 2024) causing financial losses exceeding \$3 million. We present an overview of the attack process in Figure 3. Notably, the adversary transaction that triggered the attack was labeled as an MEV Transaction on Etherscan  [^1] https://etherscan.io/tx/0x8b74995d1d61d3d7547575649136b8765acb22882960f0636941c44ec7bbe146, indicating the use of private mempool services. This allowed the attacker to bypass the public mempool and invalidate any potential front-running defense attempts. This attack highlights the limitations of traditional attack prevention mechanisms that rely on detecting adversarial transactions, particularly when private transactions are used.
@@ 3.1.1. Rescue Possibility
Despite the attacker using private mempool services to conceal the adversarial transaction, there was still an opportunity to prevent the attack, based on the following facts: 1) The ConicEthPool contract’s address — the vulnerable target — was written in plaintext in the adversary contract at the time of contract deployment; 2) There was a 36-second window (3 blocks) between the deployment of the adversarial contract and the execution of the adversarial transaction; 3) The ConicEthPool contract had an access-controlled shutdownPool function, which offers the ability to perform an emergency pause to block the adversary’s attempt to call the deposit function. Thus, detecting the adversarial contract at the deployment stage, before the confirmation of the adversarial transaction, could have prevented the exploit, regardless of whether the attacker used a private mempool service or not.
@@ 3.1.2. Adversarial Patterns
While DeFi attack logic can vary based on the specific vulnerabilities being exploited, adversarial contracts often exhibit consistent behaviours due to the nature of their malicious intent and their profit motives. Our analysis of numerous adversarial contracts reveals a clustered set of patterns including the use of anonymous fund sources, closed-source code, and frequent token-related function calls. Additionally, adversarial contracts often have distinctive interaction logic with victim contracts, which differs significantly from the patterns observed in normal contracts, resulting in easily distinguishable code structures and external call flows. Based on these observations, ML techniques can be employed to effectively identify adversarial contracts — regardless of the specific attack type — by learning and predicting these common patterns.

### v6

@@ Abstract.
The exploitation of smart contract vulnerabilities in Decentralized Finance (DeFi) has resulted in financial losses exceeding $3$ billion US dollars. Existing defense mechanisms primarily focus on detecting and reacting to adversarial transactions executed by attackers that target victim contracts. However, with the emergence of private transaction pools where transactions are sent directly to miners without first appearing in public mempools, current detection tools face significant challenges in identifying attack activities effectively.
Based on the fact that most attack logic rely on deploying intermediate smart contracts as supporting components to the exploitation of victim contracts, novel detection methods have been proposed that focus on identifying these adversarial contracts instead of adversarial transactions. However, previous state-of-the-art approaches in this direction have failed to produce results satisfactory enough for real-world deployment. In this paper, we propose LookAhead, a new framework for detecting DeFi attacks via unveiling adversarial contracts. LookAhead leverages common attack patterns, code semantics and intrinsic characteristics found in adversarial smart contracts to train Machine Learning (ML)-based classifiers that can effectively distinguish adversarial contracts from benign ones and make timely predictions of different types of potential attacks. Experiments on our labeled datasets show that LookAhead achieves an F1-score as high as $0.8966$, which represents an improvement of over $44.4\%$ compared to the previous state-of-the-art solution, with a False Positive Rate (FPR) at only $0.16\%$.
Blockchain, DeFi, Smart contract, Malware detection
*Jian Liu is the corresponding author.
copyright: ccdoi: 10.1145/3729353journalyear: 2025journal: PACMSEjournalvolume: 2journalnumber: FSEarticle: FSE083publicationmonth: 7ccs: Security and privacy Intrusion detection systems
@@ 1. Introduction
Decentralized Finance (DeFi) has gained significant traction within the blockchain ecosystem over the recent years, stimulating the rise of a diverse array of DeFi applications such as: decentralized exchanges (DEXs) (Adams et al., 2021; Martinelli and Mushegian, 2019) and lending platforms (AAV, 2020; Robert and Geoffrey, 2019). The total value locked (TVL) in these DeFi protocols has exceeded \$100 billion in mid 2024 (Def, 2024). Among the various blockchains, Ethereum (Wood et al., 2014) and Ethereum Virtual Machine (EVM)-compatible chains (Arb, 2024; BNB, 2024; Opt, 2024; Pol, 2024) have formed a particularly thriving DeFi ecosystem. At the heart of Decentralized Finance are smart contracts, which are self-executing programs deployed on blockchains, enabling trustless and transparent financial interactions without the need for intermediaries. The functionalities of smart contracts are invoked by users through transactions, after which the blockchain consensus algorithms ensure predefined actions and state transitions are execute correctly and autonomously on-chain. However, just like traditional software, smart contracts are fundamentally code-based programs, which are susceptible to vulnerabilities and various other security risks. In fact, DeFi incidents resulting from smart contract attacks have caused over 3 billion US dollars in financial losses (DeF, 2024; Lea, 2024; Zhou et al., 2023).
@@
Detection Based on Adversarial Transactions. Most existing real-time defense mechanisms leverage heuristics (Qin et al., 2023b; Wu et al., 2023; Zhang et al., 2020; Zhou et al., 2020) and Machine Learning (ML) (Gai et al., 2023; Wang et al., 2022) techniques to detect and respond (e.g. via front-running) to yet-to-be-confirmed attack transactions in public mempools. The main issue with this transaction-based detection approach is that it faces significant challenges in applying to private adversarial transactions. Namely, attackers can utilize private mempool services to send transactions directly to miners, evading visibility from other participants of the blockchain network before it’s confirmed and effectively bypassing attack detection and avoid any preventive actions from taking place. Based on our empirical study of the historical DeFi incidents over the past few years, we found that out of the $142$ DeFi attacks on Ethereum collected by us, $80$ cases ($56.3\%$) involve transactions using private mempool services. Moreover, the proportion of adversarial transactions sent using private services is observed to exhibit a significant upward trend, rising from $44.4\%$ in the first half of 2022 to a noticeably higher $77.8\%$ in the latter half of 2023.
Detection Based on Adversarial Contracts. The ability for an attacker to execute private adversarial transactions severely undermines the detection capabilities of existing transaction-based solutions, rendering them largely ineffective. In contrast, some works have redirected their focus towards developing detection strategies based on identifying adversarial contracts. As an example, Forta (For, 2024) applies NLP techniques and uses a simple logistic regression model to analyze contract bytecode for malicious intents, but it fails to achieve satisfactory performance (Table 1). Beyond Forta, the only comparable work is BlockWatchDog (Yang et al., 2024), which only support detecting reentrancy attacks. We emphasize the critical importance of designing an effective solution for detecting a wide range of DeFi attacks that can remain functional even when private mempool services are employed by the attackers.
Our method. In this paper, we propose the LookAhead system, a new framework for detecting DeFi attacks via unveiling adversarial contracts. It achieves high detection effectiveness through systematic feature design and selection, the construction of a comprehensive contract dataset, and a combination of advanced classifier models. In §3.2, we describe in detail the threat model assumptions and the scope of detection followed by our proposed system.
@@
We make an empirical observation that adversarial contracts used in DeFi attacks often contain similar patterns ($>70\%$ of the adversarial contracts use funds from anonymous sources for deployment; $>98\%$ are closed-sourced, etc.) We also note that depending on the vulnerabilities targeted, attackers write contract code in specific manners. For instance, flashloan attacks rely on complex call chains and the implementation of attack logic within callback functions. To use these behaviours to our advantage for identifying adversarial contracts, we using the stacking technique to design three ML models (transformer, classifier, and a meta classifier) that, when used together, can understand the intrinsic characteristics of malicious contracts based on code semantics and common patterns. To support the models in understanding the behaviour of Solidity code, we develop a specialized tokenization method called Pruned Semantic-Control Flow Tokenization (PSCFT) that is integrated directly into a contract bytecode decompiler.
An important observation made by (Zhou et al., 2023) is that most attacks are not executed atomically within the constructor during adversarial contract deployment, which provides a rescue time frame (after an adversarial contract has been deployed and before the attacker execute the attack logic via the contract) for defenders and victims. We use $tx_{deploy}$ to denote a contract deployment transaction, and $tx_{first}$ to represent the transaction in which the attacker initiates the actual attack via the adversarial contract. In our dataset, over $60\%$ of attacks exhibit a difference between the confirmation time of $tx_{deploy}$ and $tx_{first}$ that satisfies: $t_{first}-t_{deploy}\geq 100s$. By designing an efficient feature extraction pipeline and leveraging modern ML architectures, we can produce predictions for newly created contracts on-chain in a timely manner with amortized $t^{pred}\ll 100s$. The result of our design is LookAhead, a system that is able to effectively distinguish adversarial contracts from benign ones, and make just-in-time predictions ($t_{deploy}+t^{pred}<t_{first}$) of potential attacks.
@@
To ensure the performance and generality of our ML models, we hand-picked then carefully reviewed and identified $375$ adversarial contracts used in DeFi incidents between April 2020 and July 2024 from both the Ethereum and Binance Smart Chain (BSC) to form our adversarial contract dataset. And in total $796,437$ contracts deployed on Ethereum from June 2022 to June 2024 were collected, with $210,643$ of them being used as benign samples after filtration. We extract a selection of features from the contracts capturing patterns exhibited during both their implementation and their deployment stages. Given that our dataset contains a combination of contracts from both recent and historical attacks, we believe that our feature selection will remain reasonably relevant and effective even in the rapid growing landscape of decentralized finance.
@@
We build our classifiers using multiple supervised ML algorithms and conduct a comprehensive performance evaluation. Experiment results on our labeled datasets show that our method for identifying adversarial contracts performs exceptionally well with an F1-Score produced by the KNN meta classifier reaching as high as $0.8966$, and a false positive rate of only $0.16\%$, which represents an improvement of over $44.4\%$ compared to the previous state-of-the-art solution Forta.
@@
  • We highlight the limitations in existing transaction and contract-based detection methods and propose a new framework LookAhead that achieves significantly better performance;
  • We design a smart contract bytecode lifting and analysis pipeline with an integrated protocol to extract features and generate PSCFT for Natural Language Processing (NLP) training;
  • We construct the first large-scale comprehensive dataset of adversarial (manually-labelled) and benign (methodically selected) smart contracts consisting of a set of useful contract features based on extensive empirical observations with statistical data support;
  • We build and evaluate an ML-based system consisting of two open-weights classifiers and one meta classifier model, the results are compared with previous state-of-the-art works to substantiate our system’s effectiveness and practicality.
@@ 2.1. Ethereum & Decentralized Finance (DeFi)
Blockchain is a decentralized ledger maintained over a peer-to-peer network via consensus mechanisms. Ethereum (Wood et al., 2014) is one of the most used blockchain platforms. It was the first to introduce the Ethereum Virtual Machine (EVM) that support smart contracts, and has inspired the development of a series of EVM-compatible chains. DeFi is a blockchain-based financial ecosystem powered by smart contracts to offer financial services in a more open and transparent manner.
@@
Transactions. Transactions are used to transfer Ether, invoke contracts, or create contracts. After a transaction is sent, it is first placed in the mempool of a blockchain node, waiting to be selected and finalized by block producers.
@@ 2.2. Maximal Extractable Value (MEV)
Traditionally, miners determine the order of transactions in a block based on gas prices. Since the mempool is publicly accessible, users have the ability to influence the placement of their transactions by adjusting the gas price, allowing for front-running (Eskandari et al., 2020; Torres et al., 2021). The profit gained from such manipulation, beyond standard rewards and fees, is known as Maximal Extractable Value (Daian et al., 2020).
Front-running. Front-running is a fundamental means used to extract MEV, where bots monitor the public mempool for target transactions and raise gas prices to execute their own first. It has also been used in transaction-based intrusion prevention systems (Qin et al., 2023a; Zhang et al., 2023) to front-run adversarial transactions, thereby preventing DeFi attacks.
Private Mempool Services. In response to MEV, Flashbots’ MEV-auction has been widely adopted (Fla, 2024; Weintraub et al., 2022). This solution provides private channels for transactions to be submitted directly from the users to block producers without being broadcasted to the network, ensuring that transactions remain private until they are finalized, preventing them from being monitored or intercepted by others. However, these services have increasingly been abused by attackers to conceal adversarial transactions, effectively evading mainstream transaction-based detection methods.
@@ 3.1. Motivating Example
On July 21, 2023, DeFi protocol Conic Finance suffered a major exploit (DLNews, 2024) causing financial losses exceeding \$3 million. We present an overview of the attack process in Figure 1. Notably, the adversary transaction that triggered the attack was labeled as an MEV Transaction on Etherscan  [^1] https://etherscan.io/tx/0x8b74995d1d61d3d7547575649136b8765acb22882960f0636941c44ec7bbe146, indicating the use of private mempool services. This allowed the attacker to bypass the public mempool and invalidate any potential front-running defense attempts, highlighting the limitations of transaction-based prevention mechanisms that rely on detecting adversarial transactions.
@@ 3.1.1. Rescue Possibility
Despite the attacker using private mempool services to conceal the adversarial transaction, there was still an opportunity to prevent the attack, based on the following facts: 1) The ConicEthPool contract’s address — the vulnerable target — was written in plaintext in the adversarial contract at deployment; 2) There was a 36-second window between the deployment of the adversarial contract and the execution of the adversarial transaction; 3) The ConicEthPool contract had a shutdownPool function, which offers the ability to perform an emergency pause to block the adversary’s attempt to call the deposit function. Thus, detecting the adversarial contract at the deployment stage, before the confirmation of the adversarial transaction, could have prevented the exploit, regardless of whether the attacker used a private mempool service or not.
@@ 3.1.2. Adversarial Patterns
While DeFi attack logic can vary based on the specific vulnerabilities being exploited, adversarial contracts often exhibit consistent behaviours due to the nature of their malicious intent and their profit motives. Our analysis of numerous adversarial contracts reveals a clustered set of patterns including the use of anonymous fund sources, closed-source code, and frequent token-related function calls. Additionally, adversarial contracts often have distinctive interaction logic with victim contracts, which differs significantly from the patterns observed in normal contracts, resulting in easily distinguishable code structures and external call flows. Based on these observations, ML techniques can be employed to effectively identify adversarial contracts (regardless of the specific attack type) by learning and predicting these common patterns.